Hello tutor, may you please help me on this problem? and maybe show the work, so i know how to do my other problem, please?
One angle of a triangle is four times as great as a second angle. The third angle measures 5 degree more than twice the second angle. Find the measures of the angles.
Thank you!

Answer by mananth(16949)

let the second angle be x
first angle will be 4x
Third angle = 2x+5
The sum of angles of a triangle = 180 deg
x+4x+2x+5=180
7x+5=180
7x=180-5
7x=175
/7
x=25
second angle = 25 deg
first angle=4x=4*25=100 deg
Third angle = 2x+5=2*25+5=55 deg